Please help me solve this equation:
a^-11 / a^-8
That is: a to the power of negative 11 divide by a to the power of negative 8.
Thank you.

Answer by stanbon(75887)

a^-11 / a^-8 = a^8/a^11 = 1/a^3
===================================
Cheers,
Stan H.